This is actually a word problem.
A coffee merchant has coffee beans that sell for $9 per pound and $12 per pound. The two types are to be mixed to create a 100lb mixture that will sell for $11.25 per pound. How much of each type of bean should be used in the mixture?

Answer by checkley71(8403)

9X+12(100-X)=100*11.25 OR 9X+1200-12X=1125 OR -3X=-75 OR X=75/3 OR X=25 LBS OF $ COFFEE THEREFORE YOU NEED 75 LBS OR $12 COFFEE TO MAKE THE PROPER MIX SELLING FOR $11.25